# Measuring quantum relative entropy with finite-size effect

### v1

@@ V Analysis on time complexity
First, we apply the Schur transform. The references [20, 28, 29] proposed its quantum algorithm. We evaluate its prevision as follows. Given a required error $\epsilon>0$ and a target unitary $V$, we say that a unitary $U$ is an $\epsilon$-approximate of $V$ when any pure state $|\psi\rangle$ satisfies
@@
When the constructed unitary is an $\epsilon$-approximate of the Schur transform, the time complexity of the construction [20, 29] is $O(nd^{3}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon))$. Since these constructions realize the approximation (41), the estimation error probability increases at most only $\epsilon^{2}$. We denote the distribution of $(\lambda,j)$ when we employ an $\epsilon$-approximate of the Schur transform by $P_{\rho,\sigma}^{(n),\epsilon}$.
@@
where $\bm{\lambda}!:=\lambda_{1}!\lambda_{2}!\cdots\lambda_{d}!$. For its detail, see [26, Eq. (2.72)]. Using [21, Theorem 2.5], we have
@@
Overall, when $n=O(d^{2})$, we use the construction of Schur transform by [20, 29] with the time complexity $O(d^{5}\mathop{\rm polylog}\nolimits(d))$ and the estimate $x_{\bm{\lambda},j}$ with the time complexity $O(n)=O(d^{2})$. That is, the construction has a total time complexity $O(d^{5}\mathop{\rm polylog}\nolimits(d))$.
VI Optimality of use of Schur sampling
The paper [6] studied the case when $\sigma_{d}$ is the completely mixed state $\rho_{\mathop{\rm mix}\nolimits}$. In this case, the paper [6] showed that the lower bound of sample complexity is $O(d^{2})$ when $d$ is increasing, Schur sampling is employed. In this case, the estimate $x_{\bm{\lambda},j}$ of our estimator depends only on $\bm{\lambda}$. So, $x_{\bm{\lambda},j}$ is given as a function of the measurement outcome of Schur sampling.
In the following, we show that the optimal estimator can be given with the use of Schur sampling when $\sigma_{d}$ is the completely mixed state $\rho_{\mathop{\rm mix}\nolimits}$. That is, for any estimator POVM $M$, there exists an estimator POVM $M^{\prime}$ based on Schur sampling such that the performance of $M^{\prime}$ is not worse than the performance of $M$. For its precise statement, we introduce the following minimax criterion. When the true state is $\rho$ and the estimator POVM is $M$, we denote the distribution of the estimate $\hat{x}$ of the value $D(\rho\|\rho_{\mathop{\rm mix}\nolimits})$ by $P_{\rho;M}$. We focus on the following quantities
$R(\rho,M,\epsilon):=\max_{g\in\mathop{\rm SU}(d)}P_{g\rho g^{\dagger};M}\{\hat{x}||x-D(\rho\|\rho_{\mathop{\rm mix}\nolimits})|>\epsilon\}.$ (52)
We denote the projection to ${\cal U}_{\bm{\lambda}}\otimes{\cal V}_{\bm{\lambda}}$ by $E_{\bm{\lambda}}$. Then, we have the following lemma.
Lemma 7.
For any estimator POVM $M=\{(M_{i},\hat{x}_{i})\}_{i}$, there exists an estimator POVM $M^{\prime}=\{(M^{\prime}_{i},\hat{x}_{i})\}_{i}$ such that each operator $M^{\prime}_{i}$ is written as the form
$M^{\prime}_{i}=\sum_{\bm{\lambda}}c_{i|\bm{\lambda}}E_{\bm{\lambda}},$ (53)
and
$R(\rho,M,\epsilon)\geq R(\rho,M^{\prime},\epsilon).$ (54)
Here, $\hat{x}_{i}$ expresses the estimate when a measurement outcome $i$ is observed.
Eq. (53) means that the estimator $M^{\prime}$ is given by using Schur sampling. Eq. (54) means that the estimator $M^{\prime}$ is better than the estimator $M$. Hence, if we do not care about the time complexity, it is allowed that our estimator is a function of Schur sampling including a randomized function of Schur sampling. Hence, combining the result by [6], we find that the lower bound of sample complexity is $O(d^{2})$ when $d$ is increasing and $\sigma$ is the completely mixed state.
Proof.
To prove this lemma, we employ the method of Holevo’s covariant measurement [30, Chapter 4], [24, Chapter 4]. We denote the invariant probability measure on $\mathop{\rm SU}(d)$ by $\mu$. Since $M^{\prime}_{i}:=\int_{\mathop{\rm SU}(d)}(g^{\otimes n})^{\dagger}M_{i}g^{\otimes n}\mu(dg)$ is invariant for the representation $g^{\otimes n}$, it has the form $\sum_{\bm{\lambda}}c_{i|\bm{\lambda}}E_{\bm{\lambda}}$. We set the set $B_{\epsilon}:=\{i||x-D(\rho\|\rho_{\mathop{\rm mix}\nolimits})|>\epsilon\}$. Then, we have
$R(\rho,M,\epsilon)=\max_{g\in\mathop{\rm SU}(d)}\sum_{i\in B_{\epsilon}}\mathrm{Tr}g^{\otimes n}\rho^{\otimes n}(g^{\otimes n})^{\dagger}M_{i}$
$\geq$ $\int_{\mathop{\rm SU}(d)}\sum_{i\in B_{\epsilon}}\mathrm{Tr}g^{\otimes n}\rho^{\otimes n}(g^{\otimes n})^{\dagger}M_{i}\mu(dg)$
$=$ $\sum_{i\in B_{\epsilon}}\mathrm{Tr}\rho^{\otimes n}\Big{(}\int_{\mathop{\rm SU}(d)}(g^{\otimes n})^{\dagger}M_{i}g^{\otimes n}\mu(dg)\Big{)}$
$=$ $\sum_{i\in B_{\epsilon}}\mathrm{Tr}\rho^{\otimes n}M^{\prime}_{i}=R(\rho,M^{\prime},\epsilon).$ (55)
VII Discussion
@@
We have studied the estimation of quantum relative entropy $D(\rho\|\sigma)$ when the second argument $\sigma$ is known. Our estimator has a good performance in the following two senses. First, when the dimension $d$ is fixed, our estimator attains Cramér-Rao type bound asymptotically. Second, when the dimension $d$ increases, our estimator has sample complexity $O(d^{2})$ under natural assumptions. This sample complexity is optimal when $\sigma$ is the completely mixed state. Third, when $d$ is large, although the calculation of our estimator is difficult, we have an alternative estimator whose time complexity is $O(d^{5}\mathop{\rm polylog}\nolimits(d))$. The difference between this alternative estimator is our original estimator is zero asymptotically. In the above way, our estimator works well unifiedly, which is a unique feature and did not appear in existing studies [6, 7, 14].
@@ Acknowledgement
MH is supported in part by the National Natural Science Foundation of China (Grant No. 62171212). The author is grateful for Professor Kun Fang to a helpful discussion.
@@ Appendix A Evaluation of mean square error: Proof of Theorem 2
$=$ $\frac{1}{n^{2}}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{1}{n^{2}}(\log d_{n})^{2},$ (58)
@@ Appendix B Local asymptotic normality: Proof of Theorem 3
$=$ $\frac{1}{n}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\leq\frac{1}{n}(\log d_{n})^{2}$
@@ C.1 Preparation
$\rho^{\prime}\leq d_{n}\Gamma_{\sigma,n}(\rho^{\prime}).$ (63)
@@
$d_{n}^{\alpha-1}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
$=$ $\Big{(}d_{n}\Gamma_{\sigma,n}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}^{\alpha-1}$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
@@
$=$ $d_{n}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $d_{n}^{1-\alpha}(\mathrm{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n},$ (65)
@@ C.2 Proof of Theorem 4
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $d_{n}^{\alpha}(\mathrm{Tr}(\sigma^{\frac{\alpha}{2(1-\alpha)}}\rho\sigma^{\frac{\alpha}{2(1-\alpha)}})^{1-\alpha}e^{\alpha R})^{n}$
$=$ $d_{n}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)}.$ (66)
@@
$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n}e^{-nr}$
$=$ $\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n}e^{-nr}$
$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\mathrm{Tr}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n}e^{-nr}$
$=$ $(\mathrm{Tr}(\sigma^{-\frac{\alpha}{2(1+\alpha)}}\rho\sigma^{-\frac{\alpha}{2(1+\alpha)}})^{1+\alpha}e^{-\alpha(R-r)})^{n}+d_{n}e^{-nr}$
$=$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n}e^{-nr}.$ (67)

### v2

@@ V Analysis on time complexity
First, we apply the Schur transform $U_{Schur}$, which realizes a basis change to block-diagonalize $U^{\otimes n}$ for an arbitrary unitary $U$. The references [26, 36, 37, 38] proposed its quantum algorithm. We evaluate its prevision as follows. Given a required error $\epsilon>0$ and a target unitary $V$, we say that a unitary $U$ is an $\epsilon$-approximate of $V$ when any pure state $|\psi\rangle$ satisfies
@@
When the constructed unitary is an $\epsilon$-approximate of the Schur transform $U_{Schur}$, the time complexity of the construction [37] is $O(nd^{4}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon))$. Since these constructions realize the approximation (47), the estimation error probability increases at most only $\epsilon^{2}$. We denote the distribution of $(\lambda,j)$ when we employ an $\epsilon$-approximate of the Schur transform $U_{Schur}$ by $P_{\rho,\sigma}^{(n),\epsilon}$.
@@
where $\bm{\lambda}!:=\lambda_{1}!\lambda_{2}!\cdots\lambda_{d}!$. For its detail, see [33, Eq. (2.72)]. Using [28, Theorem 2.5], we have
@@
Overall, when $n=O(d^{2})$, we use the construction of Schur transform $U_{Schur}$ by [37] with the time complexity $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$ and the estimate $x_{\bm{\lambda},j}$ with the time complexity $O(n)=O(d^{2})$. That is, the total construction has a total time complexity $O(d^{6}\mathop{\rm polylog}\nolimits(d))$ to realize the condition (57) with a fixed value $\epsilon>0$.
Remark 7.
Initially, the paper [26] gives a polynomial algorithm for Schur transform $U_{Schur}$. However, it did not give a precise evaluation of its time complexity. Although the paper [37] writes that the protocol by the paper [26] has time complexity $O\left(nd^{3}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon)\right)$ in its Table 1, the paper [37] does not give the derivation of this evaluation because its main focus is the implementation of the mixed Schur transform $U_{MixedSchur}$, which realizes a basis change to block-diagonalize $U^{\otimes n}\otimes\overline{U}^{\otimes m}$ for an arbitrary unitary $U$. In fact, the paper [37] proved that its time complexity is $O\left((n+m)d^{4}\mathop{\rm polylog}\nolimits(n,m,d,1/\epsilon)\right)$, and the recent paper [38] also proposed the same algorithm as the algorithm by [37]. In this paper, to make a safer statement, we employ the statement that the time complexity of the unitary $U_{MixedSchur}$ is $O\left((n+m)d^{4}\mathop{\rm polylog}\nolimits(n,m,d,1/\epsilon)\right)$. Its special case with $m=0$ corresponds to $U_{Schur}$ and covers our case with the time complexity $O\left(nd^{4}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon)\right)$.
Remark 8.
The paper [7] also mentions that there exists an algorithm for Schur transform $U_{Schur}$ whose time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$ when $n=O(d^{2})$ [39]. First, they employ the algorithm for weak Schur sampling given in [40, Section 4.2.2]. Although this method employs the quantum Fourier transform the permutation group $\mathfrak{S}_{n}$, they proposed to use the algorithm by [41] for this part. The obtained algorithm has time complexity $O(n^{3}\mathop{\rm polylog}\nolimits(n,d,1/\epsilon))$. When $n=d^{2}$, it equals $O(d^{6}\mathop{\rm polylog}\nolimits(d,1/\epsilon))$. This implementation of weak Schur sampling was mentioned earlier in [42, Section 1.1.2].
VI Discussion
@@
We have studied the estimation of quantum relative entropy $D(\rho\|\sigma)$ when the second argument $\sigma$ is known. Our estimator has a good performance in the following two senses. First, when the dimension $d$ is fixed, our estimator attains Cramér-Rao type bound asymptotically. Second, when the dimension $d$ increases, our estimator has sample complexity $O(d^{2})$ under natural assumptions, which greatly improves the sample complexity by the simple application of full tomography [27]. Third, when $d$ is large, although the calculation of our estimator is difficult, we have an alternative estimator whose time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits(d))$. The difference between this alternative estimator and our original estimator approaches zero asymptotically. In the above way, our estimator works well unifiedly, which is a unique feature and did not appear in existing studies [6, 7, 16].
@@ Acknowledgement
The author is supported in part by the National Natural Science Foundation of China (Grant No. 62171212). The author is grateful for Professor Kun Fang to a helpful discussion. Also, the author would like to thank Mr Dmitry Grinko and Dr. Qisheng Wang for helpful discussions on the time complexity of Schur transform. In addition, the author would like to thank Prof. Jun Suzuki and Dr. Francesco Albarelli for helpful discussions on the Cramér-Rao type bound. Also, the author would like to thank Prof. Macro Tomamichel and Dr. Hu Yanglin for helpful discussions for the relation with full tomography and informing the references [27, 34].
@@ Appendix A Evaluation of mean square error: Proof of Theorem 2
$=$ $\frac{1}{n^{2}}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{1}{n^{2}}(\log d_{n,d})^{2},$ (60)
@@ Appendix B Local asymptotic normality: Proof of Theorem 3
$=$ $\frac{1}{n}\mathrm{Tr}\rho_{n}(\log\rho_{n})^{2}\leq\frac{1}{n}(\log d_{n,d})^{2}$
@@ C.1 Preparation
$\rho^{\prime}\leq d_{n,d}\Gamma_{\sigma,n}(\rho^{\prime}).$ (65)
@@
$d_{n,d}^{\alpha-1}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Gamma_{\sigma,n}(\rho^{\otimes n})(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha-1}$
$=$ $\Big{(}d_{n,d}\Gamma_{\sigma,n}\Big{(}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\Big{)}\Big{)}^{\alpha-1}$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $d_{n,d}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})$
@@
$=$ $d_{n,d}^{1-\alpha}\mathrm{Tr}((\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}}\rho^{\otimes n}(\sigma^{\otimes n})^{\frac{1-\alpha}{2\alpha}})^{\alpha}$
$=$ $d_{n,d}^{1-\alpha}(\mathrm{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha})^{n},$ (67)
@@ C.2 Proof of Theorem 4
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $d_{n,d}^{\alpha}(\mathrm{Tr}(\sigma^{\frac{\alpha}{2(1-\alpha)}}\rho\sigma^{\frac{\alpha}{2(1-\alpha)}})^{1-\alpha}e^{\alpha R})^{n}$
$=$ $d_{n,d}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)}.$ (68)
@@
$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n,d}e^{-nr}$
$=$ $\mathrm{Tr}\Gamma_{\sigma,n}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n,d}e^{-nr}$
$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\mathrm{Tr}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n,d}e^{-nr}$
$=$ $(\mathrm{Tr}(\sigma^{-\frac{\alpha}{2(1+\alpha)}}\rho\sigma^{-\frac{\alpha}{2(1+\alpha)}})^{1+\alpha}e^{-\alpha(R-r)})^{n}+d_{n,d}e^{-nr}$
$=$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n,d}e^{-nr}.$ (69)